OpenStudy (anonymous):

has anyone read the art of computer programming? the monograph of classical computer science? how hard is the math in those books?

OpenStudy (anonymous):

don't tell me you've worked through the second book! :O

OpenStudy (anonymous):

my physics teacher told me the thinner the american book, the harder it is! //Sorry, I was away from the keyboard.

OpenStudy (anonymous):

It's 4 books now, each about 700 pages

OpenStudy (anonymous):

here is an average question in the book (not joking): if\(f(x) = O(x^r)\) as \(x \rightarrow \infty\), show that \[\int_0^{f(x)}{e^{-u}(1+\frac{u}{x})^x}du\]\[= \int_0^{f(x)}\exp(\frac{-u^2}{2x}+\frac{u^3}{3x^2}-...+\frac{(-1)^{m-1}u^m}{mx^{m-1}})du + O(x^{-s})\]

OpenStudy (anonymous):

whoa what is that, i thought it was the art of 'programming' not art of mathematics

OpenStudy (anonymous):

i have heard about the book before but i didn't knew it has such type of math, i was thinking about reading the book this summer(2012) looks like i will have postpone it

OpenStudy (anonymous):

yeah the guy forgot to change the title of the book.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

if you can follow it, it contains lots of great algorithms in it.

OpenStudy (anonymous):

the hard part is not really the math; he asks you to program the math after each section .